### v1

# In-context Prompt Learning for Test-time Vision Recognition with Frozen Vision-language Model

Abstract
Existing pre-trained vision-language models, e.g., CLIP, have demonstrated impressive zero-shot generalization capabilities in various downstream tasks. However, the performance of these models will degrade significantly when test inputs present different distributions. To this end, we explore the concept of test-time prompt tuning (TTPT), which enables the adaptation of the CLIP model to novel downstream tasks through only one step of optimization on an unsupervised objective that involves the test sample. Motivated by in-context learning within field of natural language processing (NLP), we propose In-Context Prompt Learning (InCPL) for test-time visual recognition task. InCPL involves associating a new test sample with very few or even just one labeled example as its in-context prompt. As a result, it can reliably estimate a label for the test sample, thereby facilitating the model adaptation process. InCPL first employs a token net to represent language descriptions as visual prompts that the vision encoder of a CLIP model can comprehend. Paired with in-context examples, we further propose a context-aware unsupervised loss to optimize test sample-aware visual prompts. This optimization allows a pre-trained, frozen CLIP model to be adapted to a test sample from any task using its learned adaptive prompt. Our method has demonstrated superior performance and achieved state-of-the-art results across various downstream datasets.
1 Introduction
Recent advances in vision-language pre-training, such as CLIP [27], have shown a promising direction for developing foundation models for downstream tasks [3]. These foundation models are trained on extensive web-scale data, such as 400 million text-image pairs in the case of CLIP, to align language and vision modalities. When transferring to downstream tasks, the pre-trained model needs to be fine-tuned on a few labeled images for the task domain. Fully fine-tuning a large pre-trained model [3, 19, 6] for each downstream task faces two major challenges in real-world applications. One practical concern is the storage and distribution issue, as maintaining a separate model copy for each task is costly and inflexible, particularly with an expanding number of downstream tasks. Another one is that fully fine-tuning has large probability to destroy initial knowledge provided by the large-scale pre-trained model, increasing the risk of model overfitting.
Unlike the fully fine-tuning to expensively update all model parameters for each downstream task, prompt tuning methods [52, 51] prepend the inputs with learnable parameters that steers the foundation model towards generating the desired outputs. Recent works [52, 51, 2, 15] use training data to tune prompts where the embeddings of prompt are extracted from the model input and are differentiable with respect to the loss function. However, the learned prompts are largely limited to training data, and their performance tend to degrade significantly when the test samples are drawn from a different distribution. Test-time prompt tuning (TTPT) offers a viable solution to the distribution problem through one-step optimization on an unsupervised objective that involves only unlabelled test samples. TTPT flexibly addresses the distribution shift by prompt tuning in a zero-shot manner. For example, TPT [32] exploits the pre-trained vision-language foundation models as better zero-shot learners by maximizing the consistency of predictions across augmented versions of each test sample. However, the learned prompt via existing TTPT methods might be noisy and irrelevant to task-specific distribution. This raises the question: How can we effectively tailor a pre-trained visual-language model (e.g., CLIP) to a new task by judiciously incorporating domain-specific information during testing?
Recent works in natural language processing (NLP) have shown that in-context learning (ICL)  [^1] Without the need to fine-tune any model parameters for downstream tasks, in-context learning involves adding domain-specific input-output pairs, referred to as in-context examples, to a test example. This prompts the model to learn relevant patterns for the test sample, enabling it to perform multiple tasks seamlessly. [5] are effective in adapting foundation models for downstream tasks including knowledge retrieval [38, 41] and multi-choice tasks [24]. Leveraging In-Context Learning (ICL), large language models can tackle novel tasks solely through inference, conditioning on a limited set of input-label pairs to make predictions for new inputs. However, visual model has yet achieve this in-context learning ability. Existing works [36, 20] demonstrates that optimizing an unsupervised test-time objective, involving test samples, can enhance model performance on the target data. Inspired by this, we introduce an innovative approach for vision recognition task during test time without fine-tuning. Our approach prompts the model to acquire domain-specific context information for a test sample using just a few in-context examples.
Figure 1: Illustration of the proposed in-context prompt learning (InCPL) framework. Our approach focuses on identifying unknown objects within a contextual setting using learned prompts, all without the need for extensive fine-tuning of a large pre-trained model. See texts for details.
In this paper, we propose visual In-Context Prompt Learning (InCPL), a method that empowers a pre-trained vision-language model, such as the CLIP model, to leverage in-context examples. Our approach maintains the frozen weights of the vision-language model while enabling gradient back-propagation through it. This allows us to dynamically fine-tune the visual prompt using in-context examples and an unlabeled test sample. To achieve this, we undertake the following steps, as illustrated in Figure 1: (i) We employ a token network to convert textual descriptions into a vision prompt that the vision encoder can comprehend (Figure 1 (a)). (ii) We construct a test sample coupled with in-context candidates to harness the capabilities of the pre-trained large model (Figure 1 (b)). (iii) We design an effective learning strategy to seamlessly integrate the visual prompt with the text prompt (Figure 1 (c)). Through these steps, InCPL enables the CLIP model to adapt to new tasks by leveraging in-context information, all while keeping the model’s core parameters fixed. By harnessing the power of a pre-trained vision-language model, InCPL inherits the robust zero-shot learning capabilities of the CLIP model for downstream tasks on which it was not initially trained. During testing, InCPL seamlessly adapts to new tasks with the assistance of just a few in-context examples and a test sample. Finally, we conduct our experiments on diverse image recognition datasets and evaluate the transfer performance with pre-trained models. Our InCPL outperforms previous methods by a large margin, especially on fine-grained datasets. We also provide comprehensive ablation study and qualitative analysis to understand the effectiveness of InCPL.
Our main contributions can be summarized as follows: 1) We introduce InCPL, a straightforward yet highly effective method for enhancing the CLIP model with in-context examples. This marks the first application of visual in-context prompt learning across a range of downstream tasks. 2) We delve into the effective representation of language descriptions as visual prompt initialization. This involves the utilization of in-context examples and query samples to harness the potential of the vision encoder within the CLIP model. Additionally, we design a learning strategy for visual demonstrations that seamlessly integrate visual prompts with text prompts. 3) Our InCPL approach is rigorously evaluated through extensive experiments, demonstrating superior performance and achieving state-of-the-art (SOTA) results across a diverse set of downstream datasets.
2 Related work
Prompt Learning in Vision-Language models. Foundational vision-language (V-L) models [27, 14, 49] exploit both visual and textual modalities to encode multi-modal representations. These models are pre-trained a large amount of image-text pairs available online in a self-supervised manner. During pre-training stage, a contrastive loss function is used to pull together the features of paired images and texts while pushing away the unpaired image-text features. V-L models like CLIP [27], FILIP [47] and Florence [48] have demonstrated impressive zero- and few-shot generalization capabilities in a wide range of downstream tasks. Inspired by prompt learning in natural language process (NLP), recent works have proposed to adapt V-L models to downstream task by learning the prompt tokens. CoOp [52] prepends a category name with a prompt “a photo of a” (e.g., “a photo of a cat”), and optimizes continuous set of prompt vectors at its language branch. CoCoOp [51] further makes text prompt conditioned on each input instance (image) and thus enable dynamic text prompt to adapt to each instance. As a visual prompt technique, VP [2, 15] introduces additional random noises as task-specific learnable parameters into input image space. Their prompts are exclusively concentrated within the visual branch. Afterwards, MaPLe [16] extend CoOp [52] and VP [2] to effectively leveraging two kinds of prompts to ensure synergy between vision-language modalities during training stage. However, the learned prompts are limited to the training data distribution and their performances significantly degrade when the model is tested on inputs drawn from a different distribution. To fully delve into the utilization of visual prompts for adapting V-L models, we explore a novel approach based on in-context learning [5] to strongly encourage V-L model adapt to new tasks quickly by using only few in-context examples.
Test-time Adaptation Methods. Test-time adaptation (TTA) [10] focus on challenging scenarios with the presence of only a source model and unlabeled target data are available. This work primarily attempts to address the distribution shift by designing effective test-time objective about test sample. Two similar paradigms are test-time training (TTT)  [36, 20] and domain adaptation (DA) [29, 40]. TTT adds a self-supervised proxy task, such as recognizing rotations of an image, to training objective during training stage and computes an optimization objective at test time, while DA requires the utilization of both source and target data for training with a cross-domain loss function. Without source training data, TTA [31] does not require joint training across losses (TTT) or domain adaptation (DA). The above settings usually needs to fine-tune specific model parameters, which may discard valuable knowledge provided by the pre-trained model and increase the risk of model overfitting. TPT [32] is the first test-time prompt tuning work to enhance the zero-shot generalization of CLIP model by learning adaptive prompts for each test sample at test time. However, TPT focus only on unlabeled test sample, which leads to the prompted (retrieved) knowledge features from CLIP model might deviate from their original meanings within the target distribution.
In-context Learning. In-context learning (ICL) defined by GPT3 [5] is a new paradigm, where autoregressive language model can perform on-the-fly computational reasoning on unseen tasks [23] given prompts and examples serving as context. Flamingo [1] extends the concept of ICL to images and videos by employing language as instruction. [43] conditions the input on a pair of input-output images from the same task to indicate which task to perform. [50] further explore the impact of incontext examples for large vision models. These works  [43, 50] are developed for visual dense prediction but is not applicable to vision and language (V-L) model. Our work follows them but studies visual in-context learning from a different dimension: how to learn prompt for test-time V-L model adaptation with test sample and in-context examples.
3 Method
3.1 Preliminaries: test-time prompt learning
CLIP Models. In this paper, we focus on generalizing the vision-language pre-trained model (CLIP) [27] to downstream tasks in zero-shot manner while keeping the model parameters frozen. Without supervision from human labels, CLIP can directly learn associated visual-text representations from a large amount of image-text pairs. Specifically, CLIP trains a vision encoder $g_{I}(\cdot)$ and a text encoder $g_{T}(\cdot)$ in contrastive learning to align the embeddings and then match image-text pairs. During testing, zero-shot recognition is accomplished by comparing image features with the class names synthesized by the text encoder.
Test Time Prompt learning. With pre-trained CLIP model, test time prompt learning approaches [32] append learnable prompt tokens at the text encoder during test time, e.g., “a photo of a” is added to every class names to form “a photo of a {class}”. This provides the model very helpful context information about the downstream task. Formally, class label is wrapped within a language prompt, which can be formulated as $\{\bm{t}_{SOS},\bm{P_{\text{{t}}}},\bm{c},\bm{t}_{EOS}\}$. Here, $\bm{c}$ represents the class label, and $\bm{P_{\text{{t}}}}=\{\bm{P_{l}}\}_{l=1}^{4}$ denotes the learnable prompt initialized with word embeddings corresponding to “a photo of a”. $\bm{t}_{SOS}$ and $\bm{t}_{EOS}$ are the learnable start and end token embeddings.
Table 1: Characteristics of in-context test-time adaptation (ITTA) and other adaptations.
Setting Adaptation Available Data Loss
Train Test Source Target In-context Train loss Test loss
Fine-tuning $\times$ - $x_{t},y_{t}$ $L(x_{t},y_{t})$
Domain Adaptation $\times$ $x_{s},y_{s}$ $x_{t}$ $L(x_{s},y_{s})+L(x_{s},x_{t})$
Domain Generalization $\times$ $x_{s},y_{s}$ $L(x_{s},y_{s})$
Test-Time Training $x_{s},y_{s}$ $x_{t}$ $L(x_{s},y_{s})+L(x_{s})$ $L(x_{t})$
Fully Test-Time Adaptation $\times$ - $x_{t}$ $L(x_{t})$
ITTA $\times$ - $x_{t}$ $x_{i},y_{i}$ $L(x_{t})+L(x_{i},y_{i})$
Problem Definition. Given a pre-trained model $f_{\theta_{0}}$ with parameters $\theta_{0}$ and prompt $\bm{P}$, the proposed In-context Test-Time Adaptation (ITTA) aims to adapt $f_{\theta_{0}}$ with only few input-label pairs (known as in-context examples) $(x_{i},y_{i})$ and test samples $x_{t}$ to downstream task. During testing, ITTA fine-tunes a learnable prompt using an optimization objective. This process enhances the model’s generalizability for downstream tasks where test samples lack labels, but in-context examples are accessible. The candidate in-context example set is formed by selecting one image-label pair from each category, followed by random sampling of several pairs from this set to create in-context examples. For a more comprehensive discussion on the selection of examples, we refer to the Section 6.2 in the Appendix. Without requiring source data $(x_{s},y_{s})$, ITTA does not require joint training across domains (domain adaptation) or losses (test time training). In contrast to test-time adaptation, ITTA can provide additional in-context examples as domain-specific context information for target distribution. The difference of ITTA and other adaptations are shown in Table 1.
Figure 2: The illustration of token net.
3.2 Visual prompt input with images
The visual prompt approach enables the learning of image perturbations (such as noisy boxes or rectangles) for a frozen model. This allows the model, when prompted with these perturbations, to perform a new task. As depicted in Figure 3 (a) and (b), recent studies [2, 15] have introduced task-specific learnable parameters as visual prompts. These prompts are prepended into the input image space and are learned during the fine-tuning stage. However, these parameters are typically initialized with random noise. We contend that a general text description (e.g., a photo caption) provides richer visual context information compared to random noise.
Figure 3: The distinction between our language-aware visual prompt approach and the patched, padded, and token-based visual prompt methods.
This is because linguistic words can be perceived as interpretable lexical tokens that can be easily translated into visual content comprehensible to the vision encoder. This is especially advantageous during the test-time adaptation.
A straightforward approach to implementing a visual prompt is to combine patch image embedding with learnable parameter vectors initialized with word embeddings. However, this design not only lacks a seamless connection between the vision and language modalities but also faces the challenge of differences in dimensionality between these two types of embeddings. In this regard, we propose an efficient language-to-vision translation design that has proven to be highly effective in practice. Specifically, in addition to the word context vectors, we introduce the concept of a lightweight neural network, known as Token-Net. Token-Net is responsible for generating a conditional token vector for each input image, which is then integrated with the context visual vectors. For a visual representation of this architecture, please refer to Figure 2.
Let $\mathcal{F}_{\theta}$ denote token net parameterized by $\theta$, each visual prompt $\bm{P}_{\text{{v}}}$ is obtained by projecting text words in a prompt $\bm{P}_{\text{{t}}}$ into a set of learnable vectors via language-to-vision token net, such that $\bm{P}_{\text{{v}}}=f_{\theta}(\bm{P}_{\text{{t}}})$. Given as input image $\textbf{X}\in\mathbb{R}^{C\times H\times W}$, we divide it into $M$ pathes and produce patch tokens through a projection at vision branch of CLIP model. The visual prompt for the $i$-th input image is thus conditioned on the language description. The token net is built with a linear layer, which maps inputs of dimensions $d_{l}$ to $d_{v}$. We leave the exploration of more advanced designs to future endeavors. During training, we update the context vectors $\bm{P}_{\text{{v}}}$ together with Token-Net’s parameters $\mathbf{\theta}$. This acts as a bridge between visual and language modalities, thus reducing modality discrepancy and encouraging knowledge transfer from language to vision. Unlike CoCoOP [51], which solely conditions a text prompt on each input image without the learning of visual prompts, the explicit conditioning of our visual prompt on natural language ensure the initial visual prompts within a meaningful embedding space, thereby facilitating a smoother convergence process.
Figure 4: Illustration on the proposed visual in-context prompt learning for test-time visual recognition. Each in-context example ($x_{i},y_{i}$), test sample $x_{t}$, and its prefix text are fed into a token encoder to obtain visual, prefix, and text tokens. The text tokens are translated into a visual prompt using a token network. We optimize the visual tokens using a context-aware objective: a supervised cross-entropy term $L(x_{i},y_{i},\bm{P}_{\text{{v}}})$ involving in-context examples and an unsupervised entropy minimization $L(x_{t},\bm{P}_{\text{{v}}})$ with the test sample.
3.3 In-context learning for visual recognition
In-context learning is an emerging paradigm that involves supplying labeled examples as input to a pre-trained language model. Alongside these examples, a query example is provided for which the model must predict the label based on its understanding of the task gleaned from the provided examples. Unlike conventional few-shot learning methods, in-context learning does not necessitate the updating of all weight parameters of the pre-trained model or the learning of task-specific prompts with training data for each task. Our method, InCPLL, extends this approach into the realm of visual recognition. It does so by conditioning visual prompt learning on task-specific image-label pairs while keeping the model parameters frozen. As a result, our proposed method eliminates the need for customizing models for individual tasks and generates sample-aware prompts for the foundational model. The illustration of the proposed InCPLL is shown in Figure 4.
Formally, given an unlabeled test sample $x_{t}$ and some in-context examples $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{N}$ containing $N$ image-label pairs (i.e., an image and its class name), in-context learning can be formulated as
$y_{t}=\mathcal{F}(\mathcal{D},x_{t};\Phi).$ (1)
In-context examples $\mathcal{D}$ offers contextual guidance (prompt) to the model $\mathcal{F}(\cdot;\Phi)$, enabling it to generate the optimal $y_{t}$ for $x_{t}$ without updating model parameters. As a kind of unsupervised prompt learning [50], in-context learning relies heavily on the pre-trained model and its objective may well not align with that used in pre-training. Our InCPLL append visual prompts given as $\bm{P}_{\text{{v}}}$ with the visual tokens $\{e_{cls},e_{1},e_{2},\ldots,e_{M}\}$. Hence, the visual encoder processes the following input tokens $\widetilde{X}_{p}=\{\bm{P}_{\text{{v}}},e_{cls},e_{1},e_{2},\ldots,e_{M}\}$ to produce the prompted visual feature represented as $f_{\widetilde{p}}=\mathcal{F}(X_{p};\Phi)$. Then, the instance-specific visual prompt is learned on the following objective to provide the model with the most helpful context information about on-the-fly test sample $x_{t}$.
Test-Time Objective. Given that sample’s label is not available for test-time adaptation, we propose a context-aware unsupervised loss for visual in-context prompt learning. At each time step of adaptation, we comprise a batch of test inputs with on-the-fly test sample and in-context examples. For unlabeled test sample $x_{t}$, our test-time unsupervised objective (U-obj) is to minimize the Shannon entropy [30], $L(x_{t},\bm{P_{\text{{v}}}})=-\sum_{c}p(\hat{y}^{c})\log p(\hat{y}^{c}|x_{t},\bm{P_{\text{{v}}}})$ for the model predictions $\hat{y}=f_{\theta}(x_{t})$ of class $c$. Given image-label pairs $\mathcal{D}=\{x_{i},y_{i}\}_{i=1}^{N}$, our test-time supervised objective optimize the same loss on in-context examples with a supervised loss (S-obj) $L(x_{i},y_{i},\bm{P_{\text{{v}}}})=-\sum_{c}p(y_{i})\log p(\hat{y}^{c}_{i}|x_{i},\bm{P_{\text{{v}}}})$. We take a gradient step towards optimizing the context-aware unsupervised loss $L=L(x_{t},\bm{P_{\text{{v}}}})+\sum_{i=1}^{N}L(x_{i},y_{i},\bm{P_{\text{{v}}}})$ over learnable parameters $\bm{P_{\text{{v}}}}$ that are shared across the test batch. The overall learning objective is formulated as follows:
$\bm{P_{\text{{v}}}^{*}}=arg\min_{\bm{P_{\text{{v}}}}}\left\{L(x_{t},\bm{P_{\text{{v}}}})+{\textstyle\sum_{(x_{i},y_{i})\in\mathcal{D}}\lambda L(x_{i},y_{i},\bm{P_{\text{{v}}}},\theta)}\right\}.$ (2)
Here $\lambda$ is loss weight parameter balancing the contributions of different loss components. Each of these samples only necessitates a single pass through the visual-language model, coupled with a shared prompt. Following a single-step optimization process involving the context-aware objective, which incorporates on-the-fly test samples, we assess the online performance of the proposed InCPLL method on these test samples, utilizing the learned prompt. In the end, we utilize the average model performance on the testing data as the evaluation metric.
3.4 Cyclic Learning of Visual and Language Prompts
Visual and language (V-L) prompts are two parameter-effective prompt-tuning methods. Most methods [16, 17] attempt to simultaneously adapt vision and language branch of CLIP model with learnable prompts of both modalities, which ensures mutual synergy between V-L prompts. This mutual learning strategy performs effectively when training data provides alignment information across visual and language modalities. However, this performance is limited for the target distribution during test-time adaptation, where on-the-fly test sample is unlabeled (i.e., only visual information is available). To this end, we propose a cyclic learning of visual and language prompt to tune vision and language branch of CLIP sequentially, which leverages context information obtained from one-shot examples. Specifically, we start by optimizing one prompt and subsequently utilize the knowledge gained from this step to guide the optimization of the remaining prompt. This process provides the model with richer contextual information, allowing it to thoroughly capture relevant patterns of on-the-fly test sample to align visual and language modalities. The optimization objective for the proposed cyclic learning strategy can be formulated by
$\scriptsize\mathcal{L}(x_{t},\mathcal{D},\bm{P}_{\textbf{v}},\bm{P}_{\textbf{t}},\theta)=\begin{cases}\arg\min_{\bm{P}_{\textbf{v}},\theta}\left\{L(x_{t},\bm{P}_{\textbf{v}})+\sum_{(x_{i},y_{i})\in\mathcal{D}}\lambda L(x_{i},y_{i},\bm{P}_{\textbf{v}},\bm{P}_{\textbf{t}},\theta)\right\},s=1,\\ \arg\min_{\bm{P}_{\textbf{t}}}\left\{L(x_{t},\bm{P}_{\textbf{t}})+\sum_{(x_{i},y_{i})\in\mathcal{D}}\lambda L(x_{i},y_{i},\bm{P}_{\textbf{v}},\bm{P}_{\textbf{t}},\theta)\right\},s=2,\end{cases}$ (3)
where $x_{t}$ is unlabeled test sample, $\mathcal{D}$ contains in-context examples, $\theta$ is the parameter of token net, and $s=1,2$ are the step numbers of model optimization for visual and text prompt, respectively. It is noted that the paragraphs above section 3.4 are focused on visual prompt learning, which is a single-step optimization process. In this section, we introduce text prompts to integrate visual prompts. This process is a two-step optimization process, where each type of prompt is optimized in sequence, i.e., first visual then textual prompts. It forms the ”cyclic prompt learning” concept.
4 Experiments
Datasets. Here we evaluate the transfer performance of our method on 9 fine-grained classification datasets and 4 out-of-distribution (OOD) data of ImageNet. Fine-grained includes specific species of plant or animal species (Flowers102 [25], OxfordPets [26]), food (Food101 [4]), transportation (StanfordCars [18]), scenes [46], aircraft (Aircraft [22]), textures (DTD [7]), satellite images (EuroSAT [11]), human actions (UCF101 [35]), and general objects (Caltech101 [9]). OOD datasets is from ImageNet variants, i.e. ImageNet-V2 [28], ImageNet-Sketch [42], ImageNet-A [13] and ImageNet-R [12]. For more detailed information about the datasets and implementation details, we refer readers to Section 6.1 in the appendix.
Table 2: Accuracy comparison with previous methods on fine-grained classification datasets. CoOp and CoCoOp are fine-tuned on the ImageNet dataset using 16-shot training data per category. Baseline CLIP, prompt ensemble, and TPT do not require ImageNet dataset as training data. Our method builds above TPT and further learns sample-specific prompts for test sample by using only few in-context examples. The top-1 classification accuracy is reported on each dataset.
Method Type Flower DTD Pets Cars UCF101 Caltech Food Aircraft EuroSAT SUN Average
CLIP-ViT-B/16 Zero-shot 67.44 44.27 88.25 65.48 65.13 93.35 83.65 23.67 42.01 62.59 63.58
Ensemble [32] Zero-shot 66.99 45.04 86.92 66.11 65.16 93.55 82.86 23.22 50.42 65.63 64.59
TPT [32] Zero-shot 68.98 47.75 87.79 66.87 68.04 94.16 84.67 24.78 42.44 65.50 65.10
CoOp [52] Few-shot 68.71 41.92 89.14 64.51 66.55 93.70 85.30 18.47 46.39 64.15 63.88
CoCoOp [51] Few-shot 70.85 45.45 90.46 64.90 68.44 93.79 83.97 22.29 39.23 66.89 64.63
InCPL In-context 72.27 47.58 90.62 67.54 70.26 94.69 84.62 24.99 64.52 67.93 68.50
4.1 Generalize CLIP to fine-grained classification datasets
In this experiment, we investigate the cross-dataset generalization ability of CLIP model from ImageNet to various fine-grained datasets. Experimental results are reported in table 2. It can be seen that all prompt learning approaches perform better than the zero-shot generalization of CLIP model. However, the learned few-shot prompts (i.e., CoOp [52], CoCoOp [51]) are constrained by the distribution corresponding to training data and may exhibit limited generalization beyond that, especially for fine-grained datasets. The learned zero-shot prompts (i.e., TPT [32]) catastrophically underfit to novel tasks, whereas our model successfully learns all tasks with the help of task-specific examples. As demonstrated in Table 2, our model significantly outperforms zero-shot and few-shot baselines.
Table 3: Comparison with existing methods across a range of distribution shifts on the ImageNet dataset. CoOp, CoCoOp, CLIPOOD, MaPLe and PromptSRC are trained on source domain (ImageNet) and evaluated on out-of-domain datasets (ImageNet variant distribution). Baseline CLIP, prompt ensemble, and TPT don’t require ImageNet dataset as training data.
Method Type ImageNet ImageNet-A ImageNet-V2 ImageNet-R ImageNet-S Average
Top 1 acc. Top 1 acc. Top 1 acc. Top 1 acc. Top 1 acc.
CLIP-ViT-B/16 Zero-shot 66.73 47.87 60.86 73.98 46.09 59.11
Ensemble [32] Zero-shot 68.34 49.89 61.88 77.65 48.24 61.20
TPT [32] Zero-shot 68.98 54.77 63.45 77.06 47.94 61.44
CoOp [52] Few-shot 71.51 49.71 64.20 75.21 47.99 61.72
CoCoOp [51] Few-shot 71.02 50.63 64.07 76.18 48.75 62.13
CLIPOOD [33] Few-shot 71.60 50.4 64.9 77.2
MaPLe [16] Few-shot 70.72 50.90 64.07 76.98 49.15 62.36
PromptSRC [17] Few-shot 71.27 50.90 64.35 77.8 49.55 62.77
InCPL In-context 71.62 56.51 63.87 77.63 48.08 63.54
4.2 Generalize CLIP to distribution Shift
In this experiment, we evaluate the out-of-distribution generalization of CLIP model on different variants of ImageNet. A particular in-distribution dataset is ImageNet, while we use several out of distribution datasets from ImageNet variants, i.e. ImageNet-V2 [28], ImageNet-Sketch [42], ImageNet-A [13] and ImageNet-R [12]. The performance on four variants of ImageNet with distribution shifts are summaried in Table 3. As shown in Table 3, CoOp [52], CoCoOp [51], CLIPOOD [33], MaPLe [16] and PromptSRC [17] are few-shot prompt tuning methods, which are trained on ImageNet using 16-shot training data per category. This is a type of rapid adaptation from ImageNet to ImageNet Variants, thus achieving better classification accuracies than zero-shot generalization of CLIP model. TPT [32] learn the language prompt using only the given test sample and enhances the zero-shot performance of CLIP model without any task-specific training data. Our InCPL extends TPT to visual prompt learning and learn task-specific information from in-context examples of testing data distribution. While our InCPL is not trained on ImageNet dataset, it inherits the zero-shot abilities of CLIP model and achieve on-par generalization as ImageNet trained baselines (CoOp [52], MaPLe [16], PromptSRC [17]), surpassing existing methods in terms of average accuracy.
Table 4: Study on task and instance-specific adaptations. w/o U/S/CU-obj: without unsupervised/supervised/context-aware unsupervised objective.
Method Adaptation Flower DTD Pets
Top 1 acc. Top 1 acc. Top 1 acc.
Ours w/o U-obj Task 69.39 42.43 87.84
Ours w/o S-obj Instance 67.24 24.05 87.11
Ours (CU-obj) Task & Instance 71.13 47.34 90.60
4.3 Ablation Studies
Component Analysis. In this experiment, we assess the impact of each component within our method on both task and instance adaptation. With only a limited number of in-context examples, task-specific adaptation employs them as training data to fine-tune a task-aware prompt, which can then be applied to any test sample from a new task. In contrast, sample-specific adaptation does not depend on in-context examples; instead, it learns a sample-aware prompt using only unlabeled test samples. To investigate the effect of task- and instance-specific adaptation, We create three distinct variants of our method: (1) Ours without supervised objective (w/ U-obj) does not adapt task-specific labeled examples for each task in prompt learning; (2) Ours without unsupervised objective (w/o U-obj) does not utilize on-the-fly unlabeled test sample to learn sample-specific patterns; (3) Ours with context-aware unsupervised objective (CU-obj) adopt in-context examples as domain-specific context information and employ test sample to query model to directly predict the output label.
Results are reported in Table 4. The two methods without supervised/unsupe-rvised objectives show lower performance than ours with context-aware objective. This demonstrates that incorporating task-specific and instance-specific adaptation are both beneficial for fine-grained vision recognition problems. Task-specific adaptation is a crucial component of our problem, as instance-specific adaptation is limited to tuning prompt tokens solely on unlabeled samples using unsupervised objectives. Our method learns prompt tokens directly from task-specific image-label pairs and instance-specific test sample to adapt model accordingly.
Effect on In-context Examples and Unlabeled Test Sample. In-context examples do not belong to the same categories as the unlabeled test samples. Consequently, they provide task-specific cross-modality alignment information rather than class-specific information, as is typical in traditional few-shot learning. The test samples are unlabeled and only contribute visual modality information to the CLIP model, lacking any language information.
As shown in Table 5, visual prompt learning significantly improves the zero-shot performance of CLIP compared to language prompt learning. We also evaluate the result of our InCPL when visual and language prompts are concurrently tuned, referred to as “concurrent-prompt”, which yields inferior performance compared to single visual prompt learning. Further, we consider an image-label pair from the same categories with test sample and conduct one-shot learning
Table 5: Comparison of different prompt learning strategies with in-context examples and one-shot scenarios. V-P: vision-prompt, T-P: Text-prompt, Con-P: Concurrent-prompt, Cyc-P: Cyclic-prompt. In-con: In-context examples.
Method Type Flower DTD Pets
Top 1 acc. Top 1 acc. Top 1 acc.
CLIP-ViT-B/16 0-shot 67.44 44.27 88.25
InCPL w/ V-P In-Con 71.13 47.34 90.60
InCPL w/ T-P In-Con 69.10 45.45 88.39
InCPL w/ Con-P In-Con 69.71 46.04 88.50
InCPL w/ Cyc-P In-Con 72.27 47.58 90.62
InCPL w/ V-P 1-shot 71.99 46.57 90.35
InCPL w/ T-P 1-shot 76.65 50.77 90.81
InCPL w/ Con-P 1-shot 76.09 51.60 92.04
InCPL w/ Cyc-P 1-shot 82.62 53.84 95.86
for our approach. The one-shot can provide class-specific cross-modality alignment information for unlabeled test sample. Compared to in-context learning, one-shot learning significantly improves the performance of InCPL w/ concurrent-prompt. We argue that concurrent-prompt does not explore the class-specific contextual information of one-shot sample. As shown in Table 5, the proposed cyclic prompt learning achieve the best accuracies and outperform concurrent-prompt with large margin (e.g., +6.53 for Flower dataset). This can be attributed to the disentanglement of contextual information into different modalities through cyclic learning. Here we perform a single iteration of visual and language prompt with sequential learning, and additional iterations are expected to yield improved results. We leave the exploration of iterative training for future research.
Table 6: Accuracy of different visual prompt methods.
Method Flower DTD Pets
Top 1 acc. Top 1 acc. Top 1 acc.
Patched prompt [2] 57.90 32.51 77.98
Padded prompt [15] 56.07 32.68 79.39
Token prompt 59.32 34.63 74.19
Generic-language prompt 63.30 44.74 85.20
Ours (Language-aware) 71.13 47.34 90.25
Comparison with Previous Visual Prompt Approaches. Apart from our proposed language-aware visual prompting, we also evaluate different visual prompting methods, including pixel-level patched, padded prompts [2, 15] and sequence-level token prompt with random initialization. As shown in Figure 6, we learn a single image perturbation by adding learnable patched and padded prompt to input image, which are respectively denoted as “patched prompt” and “padded prompt”. Following [15], we also prepend some learnable parameters into input sequence of transformer, denoted as “token prompt”. Table 6 shows the accuracies with different prompt designs. Compared with patched-prompt, padded-prompt, and token-prompt, our proposed visual prompt provides a robust language-aware initialization for model optimization, which not only enhances prompt-based performance but also contributes to increased training stability. To further examine the impact of parameter initialization, we experiment with generic-language prompt, which uses a generic language initialized with random context vectors. As shown in Table 6, generic-language prompt show inferior performance to our language-aware prompt but outperforms the language-unaware prompt methods (i.e., padded, patched and token prompts). This demonstrates that incorporating language modality information into visual branch is beneficial to prompt-based visual recognition.
Figure 5: Experimental results w.r.t varied number of in-context examples.
Does the number of in-context examples matter? Recall that in-context examples are sampled from the demonstration pool, which contains one sample from each category. We are interested in understanding whether the quantity of these examples impacts performance, particularly for the supervised prompt retrieval method. To investigate this, we varied the number of in-context examples from 1 to 19, resulting in a set of results illustrated in Figure 5. Firstly, it’s evident that using in-context examples significantly outperforms the “no examples” method, and the performance clearly benefits from a larger number of examples. However, interestingly, increasing the number of examples beyond a certain point (i.e., 5) starts to decrease their performance. This phenomenon may be attributed to the fact that in-context examples can be considered as a form of “training data” , and an abundance of training data may lead the model to learn prompts specific to the in-context examples rather than the unlabeled test samples. While having more examples may not necessarily be advantageous, the crucial question lies in how to strategically select and employ these examples to harness the full potential of the CLIP model. For a comprehensive exploration and discussion of this topic, please refer to the Section 6.2 in the Appendix.
Ground Truth Matters. To study the impact of correctly-paired inputs and labels in the in-context examples, referred as “ground truth input-label mapping”, we evaluate the following three methods.
• No example is a typical test-time adaptation method that does not use any labeled data. A prediction is made via optimizing unsupervised objective involving unlabeled test sample.
• Examples w/o labels is the baseline that only uses input images without the corresponding labels, which indicates the model performance without looking at the label information.
• Examples w/ gold labels are employed in a typical in-context learning method with a set of labeled examples, which indicates the model performance with looking at relevant knowledge.
• Examples w/ random labels replace all gold labels with random labels, which are randomly sampled at uniform from label space on the testing data.
• Examples w/ same labels replace all gold labels with the same label, which is consistent with the label of test sample.
• Examples w/ oracle labels replace traditional in-context examples with the oracle examples, where labels are consistent with the label of test sample.
Figure 6: Results when using no-examples, examples w/o labels, examples w/ random labels, examples w/ same labels, examples w/ gold labels and examples w/ oracle labels in fine-grained classification task.
In Fig. 6, we present the recognition accuracy values obtained with different input-label strategies applied to in-context examples. It is evident that model performance is notably sensitive to the correctness of labels. Specifically, using correct gold labels yields better results than employing random labels. Furthermore, employing the same labels as the test sample leads to a significant improvement compared to using random labels. This observation indicates that having consistent labels provides the model with instructive information about the test samples. Moreover, employing examples with oracle labels achieves the upper bound in performance and outperforms the use of examples with gold labels. We attribute this phenomenon to the increased instructive information, which arises not only from the example labels but also from the example inputs themselves. These results underscore the substantial impact of in-context examples’ labels on In-Context Learning (ICL) performance, aligning with the findings [45].
5 Concusion
In this paper, we tackle the challenge of model generalization in scenarios where the model lacks access to source training data but possesses only a very limited number of samples (e.g., just one sample for each category). Unlike few-shot learning, which utilizes only a small number of samples as training data, we treat these samples as domain-specific contextual information for each test sample. As a result, we introduce Visual In-Context Prompt Learning, a method that empowers a pre-trained vision-language model to make full use of in-context examples. Additionally, we have developed an effective language-aware prompt along with an example selection strategy and implemented a cyclic learning technique to facilitate the seamless integration of the vision prompt with the text prompt. Our experimental results across various downstream tasks consistently demonstrate the effectiveness of our approach.
Limitation. We develop a novel in-context evaluation paradigm that improve the performance of existing prompt learning algorithms in a ChatGPT-like manner. Under this paradigm, the in-context evaluator operates on test samples at test-time with unaltered model parameters, where the model only process unlabeled test sample and in-context examples served as context information, without any other samples from the same category. The proposed InCPL requires a task-specific context dataset for each individual task. This dataset creation process inherently introduces certain priors, which can potentially limit the broad applicability of our proposed methods. In future work, we will consider developing unique context datasets tailored to different tasks. This approach would enable the adaptive selection of image-label pairs to serve as in-context prompt information for various tasks.